let f (x) = x^2 + 8x and g(x) = 1-x find (f/g) (x) and (f/g) (4)
Answer by CubeyThePenguin(3113)

(f/g)(x) = (x^2 + 8x)/(1 - x)

(f/g)(4) = (4^2 + 8(4)/(-3) = -16